derive integrated rate law for the first order reaction.
1
Expert's answer
2020-04-19T15:30:01-0400

Answer:

Let's write a general equation for the first order reaction:

A "\rightarrow" products

If at time zero (t=0) the concentration of A was equal to "a" and after certain time it decreased by "x", the rate of the reaction "v" is:

"v = -\frac{\text{d}(a-x)}{\text{d}t} = \frac{\text{d}x}{\text{d}t}"

According to the law of mass action, the rate of the reaction "v" is proportional to the product of the concentrations of the reagents, raised to the powers of stoichiometric coefficients:

"v = k(a-x)" .

Combine the right sides of the equations:

"\frac{\text{d}x}{\text{d}t} = k(a-x)" .

Separate the variables:

"\frac{\text{d}x}{a-x} = k\text{d}t" .

Integrate the both sides and get a form of the integrated rate law for the first order reaction:

"\int_{0}^{x}\frac{dx}{a-x} = \int_{0}^{t}k\text{d}t"

"\text{ln}\frac{a}{a-x} = kt"

Rearrange and get other forms of the integrated rate law for the first order reaction:

"a-x = ae^{-kt}"

"x = a(1-e^{-kt})"

Rearrange to get an expression for the rate constant of the first-order reaction:

"k = \frac{1}{t}\text{ln}\frac{a}{a-x}"

Using "a-x = a\/2" at "t = \tau_{1/2}" , we can also determine the half life of the reaction:

"\tau_{1/2} = \frac{\text{ln}2}{k}"